# CryptoX : Compositional Reasoning Evaluation of Large Language Models

### v1

@@ 3.1 Main Results
We find that as the number of encoded words increases, the reasoning difficulty also increases, leading to a decline in evaluation metrics across all models. As mentioned earlier, CryptoBench forces models to conduct compositional reasoning, requiring them to decode the question before answering it. The increase in the number of encoded words corresponds to a higher likelihood of the models conducting compositional reasoning. The experimental results indicate that while most models perform poorly in compositional reasoning tasks (we analyze specific cases in the Appendix G), o1, Gemini-Exp-1206, and Claude-3.5-Sonnet still demonstrate relatively strong performance. Open-source models are generally weaker than closed-source models, where open-source models exhibit a significant accuracy gap when answering questions compared to closed-source models.
@@
Gemini-2.0-Flash-Thinking-Exp 2.76 58.97 63.53 55.17 55.24 46.72 40.51 38.93 65.92 65.59 65.51 69.97 68.31 63.86 68.57 68.93 62.5 69.86 67.49 61.97 82.49 38.33 19.05
@@
Currently, many popular benchmarks, such as MMLU, are no longer effective at capturing the performance gaps between different models. The accuracy gaps between different models are quite limited recently, which fails to fully reflect their actual abilities. To help LLM practitioners more accurately assess the CR ability of models, we propose using the AUC to evaluate the model’s CR ability. The AUC provides a more effective measure of the model’s CR ability. In Figure 3, the redder the lines, the higher the corresponding AUC values, which also indicate stronger CR abilities of the model. It can be observed that o1, Gemini-Exp-1206, and Claude-3.5-Sonnet exhibit relatively strong CR abilities. Furthermore, the variance of the models’ AUC when evaluated in our CryptoX settings is higher than that when evaluated in the vanilla settings, as illustrated in Appendix C.7.
@@ C.6.1 ROUGE-1 Score
The statistical results of decoding accuracy using ROUGE-1Lin [2004] are shown below.
@@ C.6.2 BLEU Score
The statistical results of decoding accuracy using BLEU(1-gram)Papineni et al. [2002] are shown below.
@@ C.8 Correlation With Chatbot Arena
We calculate the Spearman correlation between CryptoBench(AUC), CryptoBench(Avg), MMLU, and Chatbot Arena. The AUC and Avg based on CryptoBench yield a higher Spearman correlation(0.32) compared to MMLU(0.13), as shown in Figure 24.

### v2

@@ 3.1 Main Results
We find that as the number of encoded words increases, the reasoning difficulty also increases, leading to a decline in evaluation metrics across all models. As mentioned earlier, CryptoBench forces models to conduct compositional reasoning, requiring them to decode the question before answering it. The increase in the number of encoded words corresponds to a higher likelihood of the models conducting compositional reasoning. The experimental results indicate that while most models perform poorly in compositional reasoning tasks (we analyze specific cases in the Appendix G), o1, o3-mini, and Gemini-2.0-Flash-Thinking-Exp-01-21 still demonstrate relatively strong performance. Open-source models are generally weaker than closed-source models, where open-source models exhibit a significant accuracy gap when answering questions compared to closed-source models.
@@
StepFun-2-16k 2.35 52.81 75.75 48.48 32.46 58.55 57.88 52.8 88.81 82.18 65.51 86.57 70.42 60.7 83.93 63.57 57.5 37.59 31.45 25.7 20.2 8.33 0.68
Gemini-1.5-Pro-001 2.09 49.06 64.53 39.96 27.02 42.07 50.83 43.84 60.95 51.98 47.39 79.51 60.91 58.6 73.93 69.29 57.86 58.51 41.34 31.34 66.67 3.67 0.0
GPT-4-Turbo 2.07 52.48 70.34 38.34 21.57 43.76 36.53 21.13 81.09 66.83 55.34 85.87 69.01 56.14 80.71 57.86 47.14 62.41 50.18 45.07 98.65 9.33 4.76
@@
Currently, many popular benchmarks, such as MMLU, are no longer effective at capturing the performance gaps between different models. The accuracy gaps between different models are quite limited recently, which fails to fully reflect their actual abilities. To help LLM practitioners more accurately assess the CR ability of models, we propose using the AUC to evaluate the model’s CR ability. The AUC provides a more effective measure of the model’s CR ability. In Figure 3, the redder the lines, the higher the corresponding AUC values, which also indicate stronger CR abilities of the model. It can be observed that o1, o3-mini, and Gemini-2.0-Flash-Thinking-Exp-01-21 exhibit relatively strong CR abilities. Furthermore, the variance of the models’ AUC when evaluated in our CryptoX settings is higher than that when evaluated in the vanilla settings, as illustrated in Appendix C.7.
@@ C.6.1 ROUGE-1 Score
The statistical results of decoding accuracy using ROUGE-1[Lin, 2004] are shown below.
@@ C.6.2 BLEU Score
The statistical results of decoding accuracy using BLEU(1-gram)[Papineni et al., 2002] are shown below.
@@ C.8 Correlation With Chatbot Arena
We calculate the Spearman correlation between CryptoBench(AUC), CryptoBench(Avg), MMLU, and Chatbot Arena. The AUC and Avg based on CryptoBench yield a higher Spearman correlation(0.61, 0.57) compared to MMLU(0.19), as shown in Figure 24.